OpenStudy (anonymous):

help with test .!!!how is the graph of y=4(2)^x-3 translated from the graph of y=4(2)^x ? il give medal.!!!!!

OpenStudy (anonymous):

The graph y=4(2)^x-3 is moved three units to the right of the original figure.

jigglypuff314 (jigglypuff314):

y = a(base)^(x-h) + k "a" is now much it's stretched or shrunk h is how much left or right k is how much up or down

OpenStudy (anonymous):

ok so the answer is 3 units to the right correct??

OpenStudy (anonymous):

Yes

jigglypuff314 (jigglypuff314):

if it is y = 4(2)^(x-3) then to the right would be correct but if it is y = 4(2)^x -3 then it would not be correct

OpenStudy (anonymous):